OpenStudy (anonymous):

Use distributive property to simplify: (28x^3v over 9) divided by (14x over 3y) Answer choices: 14x^2 over 3 2x^2y^2 over 3 2x over 3

OpenStudy (ash2326):

\[\large \frac{28x^3 \times y}{9}\div \frac{14x}{3y}\] Whenever you have a division, you can convert into multiplication by inverting the divisor like here \[\large \frac{28x^3 \times y}{9}\times \frac{3y}{14x}\] Can you try now?